# A statistical mechanical view of complex economies under the green transition

### v1

@@ I Introduction
Modern developed economies have reached remarkable levels of complexity. Nowadays, technologically sophisticated products are delivered through intricate supply chains that often involve the interaction of hundreds, or even thousands, of heterogenoues actors, ranging from designers and developers, to factories and logistics companies Pichler et al. (2023); Carvalho et al. (2021).
@@
As is well known by now, such economic development has come at the price of significant changes to our climate Shukla et al. (2022). As a consequence, our economies now face the daunting task of seeking to transition to cleaner paradigms of mass production, while at the same time retaining the economic benefits delivered by past growth. Such a challenge is often broadly referred to as the ‘green transition’, i.e., the reorientation of modern economies towards low-emission and resource-efficient ways of delivering goods to the global population.
@@
From a high level perspective, the green transition amounts to an unprecedented attempt at ‘reconfiguring’ a large complex system. There is extensive debate on the pros and cons of different pathways towards greener economies Shayegh et al. (2023), due to the importance of the tradeoffs involved and the impact they may have on the lives of countless people. On the one hand, it is rather widely accepted that — on top of the obvious gains in terms of global health — the green transition will open up new economic opportunities in the medium-long term, favouring technological innovation Acemoglu et al. (2012) and leading to the creation of new jobs Lehr et al. (2012). At the same time, though, it is also understood that the transition itself will not be painless in the short run, as it will likely result in job losses and/or in significant skill mismatches in the workforce Hanna et al. (2024), as well as in the exacerbation of preexisting socio-economic inequalities Fremstad and Paul (2019).
@@
Getting a precise quantitative understanding of the above tradeoffs and their net effects on the global economy is an almost hopeless task, due to practical impossibility of collecting fine-grained data about economic production networks from different countries Pichler et al. (2023); Bacilieri et al. (2023), which often can only be inferred at best Mungo et al. (2023). Even worse, should those data be readily available, there would still be no obvious way of aggregating them into accurate predictions of the global economy’s future trajectory, due to the unprecedented nature of the course we are on.
@@
In this paper, we seek to shed some light on the above tradeoffs by leveraging approaches that are typically used in the Statistical Mechanics of disordered systems Stein and Newman (2013), which provides a natural framework to characterize the average behavior of large complex interacting systems. The key insight provided by Statistical Mechanics is that the microscopic details of a large interacting system often become irrelevant to its macroscopic statistical description, resulting in so called self-averaging behavior. Such insight has led to massive advances in our understanding of a great variety of both physical and non-physical complex systems, ranging from spin glasses Mézard et al. (1987) and quantum fields Verbaarschot and Wettig (2000) to ecosystems Allesina and Tang (2012), machine learning algorithms Carleo et al. (2019), and game theory Galla and Farmer (2013), just to name a few.
@@
In the following, we will consider a standard General Equilibrium Theory (GET) model of an economy where profit-maximizing firms interact with each other and with utility-maximizing consumers, and seek to derive its statistical properties upon averaging over all possible interactions — within certain constraints — between such actors. Broadly speaking, GET refers to the analysis of the conditions under which supply and demand match via the interaction between rational economic agents (see, e.g., Lancaster (2012)). Despite being increasingly criticized for their somewhat unrealistic assumptions Kirman (1992); Chang et al. (2013), GET models remain very much in use among leading global economic institutions. Furthermore, a handful of studies — which our work builds upon — have already demonstrated how recasting GET models in a statistical mechanical framework leads to the emergence of results, such as sharp transitions between different economic regimes, that are in qualitative agreement with empirical observations but cannot be produced via classical GET modelling De Martino and Marsili (2005, 2006); De Martino et al. (2007); Bardoscia et al. (2012); Marsili (2014); Bardoscia et al. (2017); Moran and Bouchaud (2019).
@@ II Large Random Economies
We consider an economy with $N$ firms, $C$ goods and one representative consumer, a concept often used in GET modelling to capture average/typical consumer behaviour Lancaster (2012). Each firm is endowed with a technology that allows it to produce certain goods by consuming others. This is captured by a matrix with entries $q_{n}^{c}$ that are positive (negative) when firm $n$ produces (consumes) good $c$ ($n=1,\ldots,N$; $c=1,\ldots C$). Each firm generates a certain amount of waste $\epsilon_{n}>0$, i.e.,
$\sum_{c=1}^{C}q_{n}^{c}=-\epsilon_{n}\ ,$ (1)
which effectively quantifies a firm’s efficiency as the discrepancy between the output produced by its technology and the input required to produce it. Firms operate their technologies at a scale $s_{n}>0$, with $s_{n}=0$ indicating that firm $n$ is inactive. When active, firms make a profit
@@
It can be shown Lancaster (2012); De Martino et al. (2007); Marsili (2014); Bardoscia et al. (2017) that the joint solution of the two above optimization problems, coupled with the condition in Eq. (3), is equivalent to solving
@@
Our goal in the following is to characterize the economy’s properties irrespective of its details. In order to do so, we will ‘promote’ $\bm{q}$, $\bm{\epsilon}$ and $\bm{x}_{0}$ to random variables, and take the ‘thermodynamic’ limit $N,C\rightarrow\infty$ with $\nu=N/C$ fixed. As we will see, under such a limit the economy’s main statistical properties display self-averaging behaviour and attain typical values that do not depend on the economy’s specificities.
Let us reformulate the optimization problem in Eq. (5) as
where
The above expression obviously mirrors a canonical partition function in Statistical Mechanics, and the optimization problem in Eq. (6) has been converted into a steepest descent one by the limit on the ‘inverse temperature’ $\beta\rightarrow\infty$ that selects the values of the utility function $U_{N}$ that dominate the integral above. Reasonably, we can expect the maxima of the function $U_{N}$ to grow with $N$, so that the quantity $U_{N}/N$ in Eq. (6) will be an intensive one. Our position is that, for a reasonable choice of the utility function, the maxima of $U_{N}/N$ will cease to depend on the economy’s details as it grows large, i.e.,
$\lim_{N\rightarrow\infty}\frac{1}{N}\max_{\bm{s}\geq 0}\ U_{N}(\bm{s}|\bm{q},\bm{\epsilon},\bm{x}_{0})=\lim_{N\rightarrow\infty}\frac{1}{N}\left\langle\max_{\bm{s}\geq 0}\ U_{N}(\bm{s}|\bm{q},\bm{\epsilon},\bm{x}_{0})\right\rangle_{\bm{q},\bm{\epsilon},\bm{x}_{0}}\ ,$
where $\langle\cdots\rangle_{\bm{q},\bm{\epsilon},\bm{x}_{0}}$ denotes averaging over the probability density functions describing the statistical properties of the economy’s production matrix, waste, and initial endowment.
From a practical standpoint, the above expression looks rather inconvenient, as it requires averaging over the logarithm of the partition function $Z_{N}$. Such a rather cumbersome task is usually circumvented via the so-called replica trick, which converts the average over $\log Z_{N}$ into an average over $r$ replicas of the system through the following formal identity
$\langle\log Z_{N}\rangle=\lim_{r\rightarrow 0}\frac{\langle Z_{N}^{r}\rangle-1}{r}\ ,$
where $r$ is taken to be an integer number throughout all calculations, and restored to be a real number only at the end when taking the above limit.
All in all, the optimization problem in Eq. (6) can be solved as
$\lim_{N\rightarrow\infty}\frac{1}{N}\left\langle\max_{\bm{s}\geq 0}\ U_{N}(\bm{s}|\bm{q},\bm{\epsilon},\bm{x}_{0})\right\rangle_{\bm{q},\bm{\epsilon},\bm{x}_{0}}=\lim_{N\rightarrow\infty}\lim_{\beta\rightarrow\infty}\lim_{r\rightarrow 0}\frac{1}{\beta Nr}\log\langle Z_{N}^{r}(\beta|\bm{q},\bm{\epsilon},\bm{x}_{0})\rangle_{\bm{q},\bm{\epsilon},\bm{x}_{0}}=h(\bm{\theta}^{\star})\ ,$ (7)
where the last equality comes from the fact that the replica method typically allows to identify order parameters $\bm{\theta}=(\theta_{1},\ldots,\theta_{k})$ such that
$\langle Z_{N}^{r}(\beta|\bm{q},\bm{\epsilon},\bm{x}_{0})\rangle_{\bm{q},\bm{\epsilon},\bm{x}_{0}}\approx\int\mathrm{d}\bm{\theta}\ \mathrm{e}^{\beta Nrh(\bm{\theta})}$ (8)
@@ III Solution and typical properties
In Appendix A we provide a detailed analytical solution to the optimization problem in Eq. (7) under the following assumptions:
  • As customary Lancaster (2012), we choose the utility function in Eq. (2) to be concave and monotonically increasing by choosing $u(x^{c})=\log(x^{c})$
@@
    Throughout the rest of the paper we choose $\epsilon_{0}=g\lambda$ with $g=10$ in order to effectively make sure that $\epsilon_{n}>0$, as it should be. Albeit inelegant, this assumption is a necessary compromise to retain the Gaussian form of the above probability distribution, which allows to make analytical progress. Thus, in the following, larger values of $\lambda$ will correspond to higher waste, as they lead both to a shift in the mean and to a broader variation in the distribution in Eq. (9).
  • We assume each row of the production matrix to be drawn from a Gaussian distribution with mean zero and variance $1/C$, constrained to match the constraint in Eq. (1) for fixed $\epsilon_{n}$ (drawn as explained in the above point):
@@
    where $\delta(\cdot)$ is a Dirac delta.
@@
where
$h\left(\Omega,\kappa,p,\sigma,\chi,\hat{\chi}\right)$ $=$ $\frac{\Omega\hat{\chi}}{2}-\frac{\sigma^{2}\chi}{2\nu}-\frac{\chi\left(1+\lambda\right)p^{2}}{2\nu}+\frac{p\kappa}{\nu}+\left<\max_{s\geq 0}\left[-\frac{\hat{\chi}}{2}s^{2}+\left(t\sigma-p\epsilon_{0}\right)s\right]\right>_{t}$
$+$ $\frac{1}{\nu}\left\langle\max_{s\geq 0}\left[u\left(x\right)-\frac{1}{2\chi}\left(x-x_{0}+t\sqrt{\nu\Omega}+\frac{\kappa}{1+\lambda}\right)^{2}\right]\right\rangle_{t,x_{0}}\ .$
We provide a detailed derivation of the above result in Appendix A. In the above expression, $\bm{\theta}=(\Omega,\kappa,p,\sigma,\chi,\hat{\chi})$ denote the six order parameters that characterize the solution, whose optimal values are to be determined by solving the saddle point equations
$\left(\frac{\partial h}{\partial\Omega}\right)_{\Omega=\Omega^{\star}}=\left(\frac{\partial h}{\partial\kappa}\right)_{\kappa=\kappa^{\star}}=\left(\frac{\partial h}{\partial p}\right)_{p=p^{\star}}=\left(\frac{\partial h}{\partial\sigma}\right)_{\sigma=\sigma^{\star}}=\left(\frac{\partial h}{\partial\chi}\right)_{\chi=\chi^{\star}}=\left(\frac{\partial h}{\partial\hat{\chi}}\right)_{\hat{\chi}=\hat{\chi}^{\star}}=0\ .$ (13)
The averaging with respect to $t$ in Eq. (III) simply denotes an average with respect to a standard Gaussian variable, i.e., $\langle\cdots\rangle=(2\pi)^{-1/2}\int_{-\infty}^{\infty}\mathrm{d}t(\cdots)\mathrm{e}^{-t^{2}/2}$.
@@ IV Results
In Appendix A we detail the explicit form of the saddle point equations in Eq. (13), which must be solved numerically. Their solution provides access to a number of key quantities that capture the typical behaviour of the economy as a function of its main parameters. More specifically, in the following we present our results focusing on ($i$) the ratio $\nu=N/C$ (number of available technologies per good), which we take as a proxy for the economy’s development level; ($ii$) the $\lambda$ parameter that characterizes both the average and the variation in the distribution of waste generated by different technologies (in short, the higher $\lambda$ the more the economy’s technologies generate waste, see Eq. (9)); ($iii$) the fraction $f_{0}$ of primary goods in the economy (Eq. (4)).
Figure 1: Phase diagram of the model in the ($\nu,\lambda$) plane, for different values of $f_{0}$. Colors in the heatmaps correspond to equilibrium values of the optimal average scale of production $\langle s^{\star}\rangle$.
In Fig. 1 we report the average optimal scale of production $\langle s^{\star}\rangle$ (for which we derive a closed-form expression in Appendix B, see Eq. (21)) at economic equilibrium in the $(\nu,\lambda)$ plane, both in the case of a resource-rich ($f_{0}=0.8$, left panel) and a resource-poor ($f_{0}=0.4$, right panels) economy. As it can be seen, as $\nu$ increases the economy undergoes two sharp transitions, delimiting a region in which there is economic activity ($\langle s^{\star}\rangle>0$) from two regions in which no technology is active and the economy is effectively shut down. The nature of such transitions will be explored in detail in a later Section.
@@
The plots in Fig. 2 provide some food for thought concerning an economy’s path towards waste reduction, an obviously salient point in view of most real-world economies’ efforts to curb emissions and transition towards ‘greener’ industrial paradigms. In the top-left panel we show sample horizontal ‘slices’ of the heatmaps in Fig. 1 corresponding to the average scale of production $\langle s^{\star}\rangle$ as a function of $\nu$ in four sample scenarios obtained combining the values $f_{0}=0.8$ (resource-rich economy) and $f_{0}=0.4$ (resource-poor economy) with $\lambda=0.001$ (‘green’ economy) and $\lambda=0.05$ (‘brown’ economy). Together with $\langle s^{\star}\rangle$, for the same combinations of values we also plot the average fraction of active firms/technologies $\phi=\lim_{N\rightarrow\infty}\sum_{n=1}^{N}\Theta(s_{n}^{\star})/N$ (top-right panel), the economy’s GDP $G$ (bottom-left panel), and the average consumption $\langle x^{\star}\rangle$ (bottom-right panel), see Appendix B for closed or semi-closed form expressions for such quantities.
@@
As a sanity check, we tested the validity of the analytical results we obtained from the (numerical) solution of the saddle point equations (see Eq. (13) and Appendix A) against the ‘brute force’ solution of the problem in Eq. (5) for finite $N$ via numerical optimization. Sample results for $N=150$ are presented in Fig. 3. As it can be seen, we find very good agreement between numerical simulations and theory.

### v2

@@ I Introduction
Modern developed economies have reached remarkable levels of complexity. Nowadays, technologically sophisticated products are delivered through intricate supply chains that often involve the interaction of hundreds, or even thousands, of heterogenoues actors, ranging from designers and developers, to factories and logistics companies [1, 2].
@@
As is well known by now, such economic development has come at the price of significant changes to our climate [3]. As a consequence, our economies now face the daunting task of seeking to transition to cleaner paradigms of mass production, while at the same time retaining the economic benefits delivered by past growth. Such a challenge is often broadly referred to as the ‘green transition’, i.e., the reorientation of modern economies towards low-emission and resource-efficient ways of delivering goods to the global population.
@@
From a high level perspective, the green transition amounts to an unprecedented attempt at ‘reconfiguring’ a large complex system. There is extensive debate on the pros and cons of different pathways towards greener economies [4], due to the importance of the tradeoffs involved and the impact they may have on the lives of countless people. On the one hand, it is rather widely accepted that—on top of the obvious gains in terms of global health—the green transition will open up new economic opportunities in the medium-long term, favouring technological innovation [5] and leading to the creation of new jobs [6]. At the same time, though, it is also understood that the transition itself will not be painless in the short run, as it will likely result in job losses and/or in significant skill mismatches in the workforce [7], as well as in the exacerbation of preexisting socio-economic inequalities [8].
@@
Getting a precise quantitative understanding of the above tradeoffs and their net effects on the global economy is an almost hopeless task, due to practical impossibility of collecting fine-grained data about economic production networks from different countries [1, 9], which often can only be inferred at best [10]. Even worse, should those data be readily available, there would still be no obvious way of aggregating them into accurate predictions of the global economy’s future trajectory, due to the unprecedented nature of the course we are on.
@@
In this paper, we seek to shed some light on the above tradeoffs by leveraging approaches that are typically used in the Statistical Mechanics of disordered systems [11], which provides a natural framework to characterize the average behavior of large complex interacting systems. The key insight provided by Statistical Mechanics is that the microscopic details of a large interacting system often become irrelevant to its macroscopic statistical description, resulting in so called self-averaging behavior. Such insight has led to massive advances in our understanding of a great variety of both physical and non-physical complex systems, ranging from spin glasses [12] and quantum fields [13] to ecosystems [14], machine learning algorithms [15], and game theory [16], just to name a few.
@@
In the following, we will consider a standard General Equilibrium Theory (GET) model of an economy where profit-maximizing firms interact with each other and with utility-maximizing consumers, and seek to derive its statistical properties upon averaging over all possible interactions—within certain constraints—between such actors. Broadly speaking, GET refers to the analysis of the conditions under which supply and demand match via the interaction between rational economic agents (see, e.g., [17]). Despite being increasingly criticized for their somewhat unrealistic assumptions [18, 19], GET models remain very much in use among leading global economic institutions. Furthermore, a handful of studies—which our work builds upon—have already demonstrated how recasting GET models in a statistical mechanical framework leads to the emergence of results, such as sharp transitions between different economic regimes, that are in qualitative agreement with empirical observations but cannot be produced via classical GET modelling [20, 21, 22, 23, 24, 25, 26].
@@ II Large Random Economies
We consider an economy with $N$ firms, $C$ goods and one representative consumer, a concept often used in GET modelling to capture average/typical consumer behaviour [17]. Each firm is endowed with a technology that allows it to produce a certain bundle of goods by consuming others. This is captured by a matrix with entries $q_{n}^{c}$ that are positive (negative) when firm $n$ produces (consumes) good $c$ ($n=1,\ldots,N$; $c=1,\ldots C$). Such a matrix is somewhat akin to the tables compiled in input-output (IO) analysis [27, 28]—the branch of Economics that seeks to model the interdependencies between different sectors of an economy and/or different economies. In that analogy, $q_{n}^{c}$ would quantify, e.g., the amount of goods produced by sector (country) $n$ that sector (country) $c$ consumes/imports in order to achieve its own production goals.
The matrix entries $q_{n}^{c}$ in the above framework effectively represent technical coefficients, expressing the amounts of goods produced and consumed by firms in some suitable unit measure. As such, one can choose a scale such that a firm’s total output equals one, i.e.,
$\sum_{c=1}^{C}\Theta(q_{n}^{c})=1\ ,$
where $\Theta(x)=1$ for $x>0$ (and $\Theta(x)=0$ otherwise) only selects those goods that are produced by firm $n$. With this position, we then assume that each firm generates a certain amount of waste $\epsilon_{n}>0$ per output unit, i.e.,
$\sum_{c=1}^{C}q_{n}^{c}=-\epsilon_{n}\ .$ (1)
The above quantity captures the a firm’s efficiency as the discrepancy between one unit of the output produced by its technology, and the input required to produce it. As such, in the following we will relate it to the undesired byproducts—most notably emissions—associated with production of technologically advanced products on a large scale.
Firms operate their technologies at a scale $s_{n}>0$, with $s_{n}=0$ indicating that firm $n$ is inactive. When active, firms make a profit
@@
It can be shown [17, 22, 24, 25] that the joint solution of the two above optimization problems, coupled with the condition in Eq. (3), is equivalent to solving
@@
Our goal in the following is to characterize the economy’s properties irrespective of its details. In order to do so, we will leverage ideas and methods borrowed from the Statistical Mechanics of disordered systems. Namely, we will ‘promote’ $\bm{q}$, $\bm{\epsilon}$ and $\bm{x}_{0}$ to random variables, and seek to describe the economy’s statistical properties when it grows large, i.e., when both the number of firms and goods grow to infinity at a fixed rate, $N,C\rightarrow\infty$ with $\nu=N/C$ fixed. In Statistical Mechanics, this is known as the thermodynamic limit. Under such limit—as is often the case with large interacting systems—the economy’s main statistical properties will display self-averaging behaviour and attain typical values that do not depend on the economy’s specificities. In a nutshell, this amounts to
$\lim_{N\rightarrow\infty}\frac{1}{N}\max_{\bm{s}\geq 0}\ U_{N}(\bm{s}|\bm{q},\bm{\epsilon},\bm{x}_{0})=\lim_{N\rightarrow\infty}\frac{1}{N}\left\langle\max_{\bm{s}\geq 0}\ U_{N}(\bm{s}|\bm{q},\bm{\epsilon},\bm{x}_{0})\right\rangle_{\bm{q},\bm{\epsilon},\bm{x}_{0}}\ ,$ (6)
where $\langle\cdots\rangle_{\bm{q},\bm{\epsilon},\bm{x}_{0}}$ denotes averaging over the probability density functions describing the statistical properties of the economy’s production matrix, waste, and initial endowment (which we will specify in the next Section). The above result, coupled with Eq. (5), ensures that all economies will ‘look alike’, as long as they are large enough.
In order to keep the paper’s focus on our main results, we briefly outline the procedure to solve the above optimization problem in Appendix A, which the reader well versed in the Statistical Mechanics of disordered systems can safely skip. The reader interested in a thorough and considerably more exhaustive illustration than the sketch provided in Appendix A is referred to the excellent review in [29].
@@ III Solution and typical properties
In Appendix B we provide a detailed analytical solution to the optimization problem in Eq. (6) under the following assumptions:
  • As customary [17], we choose the utility function in Eq. (2) to be concave and monotonically increasing by choosing $u(x^{c})=\log(x^{c})$
@@
    Throughout the rest of the paper we choose $\epsilon_{0}=g\lambda$ with $g=10$ in order to effectively make sure that $\epsilon_{n}>0$, as it should be. Albeit inelegant, this assumption is a necessary compromise to retain the Gaussian form of the above probability distribution, as it is often prohibitive to make any analytical progress in replica trick calculations—the analytical methodology we are using in order to average over economies, see Appendices A and B—with non-Gaussian distributions. Moreover, let us also emphasise that one can safely expect the final results not to change qualitatively as long as the pdf in Eq. (7) remains defined on an effectively positive support with variance proportional to $\lambda$. In this respect, the value of $g$ does not play any relevant role, as long as it ensures that such conditions are met.
  • We assume each row of the production matrix to be drawn from a Gaussian distribution with mean zero and variance $1/C$. This, however, would not guarantee that the constraint in Eq. (1) is met for fixed $\epsilon_{n}$ (drawn from the pdf detailed in the above point). In order to make sure of that, we enforce the constraint on each technology with a Delta function, resulting in the following pdf:
@@
    where the denominator ensures proper normalization. All in all, the above form guarantees that the vectors $\bm{q}_{n}$ are normally distributed multivariate random variables that exactly match the desired constraints.
@@
where $h$ is a function which we compute in Appendix B. Its arguments are the so called order parameters, which can be identified with the statistical mechanical approach we outline in Appendix A, and explicitly evaluated as the solution to a set of coupled non-linear equations, known as saddle point equations (see Eq. (21) in Appendix B).
@@ IV Results
In Appendix B we detail the explicit form of the saddle point equations in Eq. (21), which must be solved numerically. Their solution provides access to a number of key quantities (see Appendix C) that capture the typical behaviour of the economy as a function of its main parameters. More specifically, in the following we present our results focusing on ($i$) the ratio $\nu=N/C$ (number of available technologies per good), which we take as a proxy for the economy’s development level; ($ii$) the $\lambda$ parameter that characterizes both the average and the variation in the distribution of waste generated by different technologies (in short, the higher $\lambda$ the more the economy’s technologies generate waste, see Eq. (7)); (iiiiii) the fraction $f_{0}$ of primary goods in the economy (Eq. (4)).
Figure 1: Phase diagram of the model in the development ($\nu$) – waste ($\lambda$)
plane, for different values of the fraction of primary goods $f_{0}$. Colors in the heatmaps correspond to equilibrium values of the optimal average scale of production $\langle s^{\star}\rangle$. As it can be seen, in both plots the ($\nu,\lambda$) plane features a region characterized by positive economic output ($\langle s^{\star}\rangle>0$), which separates two regions of economic shutdown ($\langle s^{\star}\rangle=0$).
In Fig. 1 we report the average optimal scale of production $\langle s^{\star}\rangle$ (for which we derive a closed-form expression in Appendix C, see Eq. (22)) at economic equilibrium in the $(\nu,\lambda)$ plane, both in the case of a resource-rich ($f_{0}=0.8$, left panel) and a resource-poor ($f_{0}=0.4$, right panels) economy. As it can be seen, as $\nu$ increases the economy undergoes two sharp transitions, delimiting a region characterized by positive economic output ($\langle s^{\star}\rangle>0$) from two regions in which no technology is active and the economy is effectively shut down ($\langle s^{\star}\rangle=0$). The nature of such transitions will be explored in detail in a later Section.
@@
The plots in Fig. 2 provide some food for thought concerning an economy’s path towards waste reduction, an obviously salient point in view of most real-world economies’ efforts to curb emissions and transition towards ‘greener’ industrial paradigms. In the top-left panel we show sample horizontal ‘slices’ of the heatmaps in Fig. 1 corresponding to the average scale of production $\langle s^{\star}\rangle$ as a function of $\nu$ in four sample scenarios obtained combining the values $f_{0}=0.8$ (resource-rich economy) and $f_{0}=0.4$ (resource-poor economy) with $\lambda=0.001$ (‘green’ economy) and $\lambda=0.05$ (‘brown’ economy). Together with $\langle s^{\star}\rangle$, for the same combinations of values we also plot the average fraction of active firms/technologies $\phi=\lim_{N\rightarrow\infty}\sum_{n=1}^{N}\Theta(s_{n}^{\star})/N$ (top-right panel), the economy’s GDP $G$ (bottom-left panel), and the average consumption $\langle x^{\star}\rangle$ (bottom-right panel), see Appendix C for closed or semi-closed form expressions for such quantities.
@@
As a sanity check, we tested the validity of the analytical results we obtained from the (numerical) solution of the saddle point equations (see Eq. (21) and Appendix B) against the ‘brute force’ solution of the problem in Eq. (5) for finite $N$ via numerical optimization. Sample results for $N=150$ are presented in Fig. 3. As it can be seen, we find very good agreement between numerical simulations and theory.